# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v2

@@ A.3 Retrieval-Augmented Approaches
The retrieval-augmented approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) enhance LLMs by incorporating relevant documents from external corpus. These approaches represent a promising direction for addressing the knowledge limitations of LLMs, such as the need for knowledge updates (Jang et al., 2022) and long-tail knowledge (Kandpal et al., 2023). Early retrieval-augmented methods (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022) focused on the jointly training of LLMs and retrieval modules to improve their cooperation. With the evolution of general-purpose LLMs, recent studies (Ram et al., 2023; Shi et al., 2023) investigate the strategy of appending relevant documents directly to the input while keeping the LLMs frozen. Despite the effectiveness of these methods, the retrieval-augmented approaches still face challenges due to irrelevant retrieval results and incomplete knowledge coverage Yu et al. (2023); Mallen et al. (2023). These noisy retrieval results can misguide the outcomes of LLMs (Mallen et al., 2023; Yoran et al., 2023; Ren et al., 2023).
A.4 Hybrid Approaches and Knowledge Conflicts
@@
Recent works investigate to merge retrieved and generated contexts to leverage both parametric knowledge and external knowledge Abdallah and Jatowt (2023); Yu et al. (2022). These combination methods have shown improved performance over those relying solely on a single information source in fully-supervised setting. Furthermore, (Zhang et al., 2023) proposes an improved method to effectively leverage the two sources of information, especially when conflicts arise. While these works have focused on improving the efficacy of hybrid approaches, the underlying mechanisms by which LLMs process conflicting information from different types of contexts remain underexplored.
@@
Current research on knowledge conflicts in LLMs primarily focuses on two aspects: conflicts between input contexts and LLMs’ parametric memory, and conflicts among the input contexts themselves. Regarding the former, Xie et al. (2023); Chen et al. (2022) find that LLMs are highly receptive to the input contexts rather than their internal memory. Concerning conflicts within multiple input contexts, Chen et al. (2022) demonstrates that LLMs tend to rely on a few most relevant retrieved contexts. Additionally, Xie et al. (2023) reveals a confirmation bias, i.e., LLMs demonstrate a tendency to favor contexts that align with their parametric knowledge, when confronted with both supporting and opposing contexts. However, these studies are limited to analyzing context conflicts within a single type of input contexts. Our work complements these studies by considering conflicts between generated and retrieved contexts, and reveals several key factors, such as semantic completeness and text similarity.
A.5 Length Control for Generated Contexts
A.5.1 Length Distribution across LLMs
@@
In our proposed framework, we regulate the length of generated contexts by incorporating length constraint in the prompt: Generate a background context from Wikipedia to answer the given question {#question} Keep the length of the document around $n$ words We observed that GPT 4 effectively controls the output length, whereas other models struggle with this aspect. To address this issue in the latter, we employ multiple values of $n$ and select the one that best matched the retrieved context.
As the result, Figure 9 shows the length distribution of retrieved contexts and contexts generated by various LLMs. The length distribution of retrieved contexts is more concentrated due to the fact that these contexts consist of text truncated precisely at one hundred words, coupled with their titles (Karpukhin et al., 2020). Variations in the length of different retrieved articles are solely attributable to the differences in title lengths.
A.6 Length Distribution with Different Control Methods
Figure 10 illustrates the length distribution for generated contexts corresponding to three different length control methods. From the results, we can observe that the contexts generated by GenRead (Yu et al., 2022) are significantly longer compared to the retrieved contexts. This difference in length could potentially influence the model’s bias or inclination.
Table 7 illustrates the statistics of three types of generated contexts with different methods to control the length. These contexts show similar average length and similarity, with the preliminary difference in the completeness.x
Context Completeness NQ-AIR TQA-AIR Sentence Semantic Length Jaccard Length Jaccard GPT-3.5 Nature 106.0 0.187 105.0 0.398 S-Trunc 104.9 0.184 106.3 0.397 Trunc 107.4 0.187 103.6 0.397 Llama2-13b Nature 109.7 0.202 105.5 0.313 S-Trunc 105.9 0.193 104.8 0.2934 Trunc 107.4 0.196 106.3 0.2954
Table 7: Average length and similarity of generated context with different control methods. Three types of methods create context with comparable average length and similarity. The core difference lies in the completeness.
@@ A.7 Dataset Statistics
Table 8: The data quantities of the constructed subsets for different (Generator, Reader) pairs. NQ and TQA refer to the original complete test sets.
A.8 Effect of Context Order
@@
generated-first 69.9 68.2
retrieved-first 66.5 55.6
random 69.1 58.6
Table 9: The generation ratio (%) with different context order on NQ-AIR and TQA-AIR datasets. The backbone LLM is GPT 3.5.
In above experiments, retrieved and generated contexts are presented with random order. Previous studies Xie et al. (2023); Liu et al. (2023); Lu et al. (2022) have found that the model may be sensitive to the order of the input contexts. In their experiments, the input context was either all retrieved Liu et al. (2023) or all generated Xie et al. (2023). We conducted experiments to investigate whether the context order impacts the preference for the generated context. The generated and retrieved contexts are concatenated with three different orders: generated-first, retrieved-first and random. We compute the context preference with different context order respectively.
As shown in Table 9, across all context orders, LLMs consistently show a strong tendency to generated contexts (generation ratio > 0). When the retrieved context is positioned first, there is a slight reduction in the generation ratio. This reduction may result from the LLMs’ preference for generated contexts being partially offset by their bias towards the top context (Liu et al., 2023; Xie et al., 2023).
A.9 More Results on AIG Datasets
Figure 11 shows the $\mathrm{DiffGR}$ with different (reader, generator) pairs on their corresponding NQ-AIG datasets. It can be observed that LLMs show strong tendency to rely on generated contexts across various (reader, generator) pairs.
Figure 11: $\mathrm{DiffGR}$ (%) with different (reader, generator) pairs on their corresponding NQ-AIG datasets.
A.10 More Details about Confirmation Bias
A.10.1 Experiment Setting
We construct counter-memory answers and contexts based on the original AIR context-conflicting subsets, as outlined below:
@@ A.10.1 Experiment Setting
Give you a reference word, transform it into a analogous word. Reference: Missouri River Analogous Word: Mississippi River Reference: {memory answer} Analogous Word:
Furthermore, to facilitate the calculation of $\mathrm{DiffGR}$, we only retain those samples where the counter-memory answer also diverged from $a_{\phi}^{\gamma}$, the answer provided by the retrieved context. Formally, the counter-memory answer satisfies $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\varrho}$ and $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\gamma}$.
@@
Answer Consistency Checking. To verify that the counter-memory context actually supports the counter-memory answer, we retain only those instances where the predicted answer, derived exclusively from the counter-memory context, exactly match the counter-memory answer:
@@
Context NQ-AIR TQA-AIR Length Jaccard Length Jaccard Retrieved 107.5 0.111 106.3 0.184 GPT-3.5 Gen 105.9 0.194 104.9 0.385 Ctr 105.7 0.251 103.8 0.499 Llama2-13b Gen 107.4 0.197 104.3 0.319 Ctr 106.7 0.165 103.5 0.310
Table 10: Average length and Jaccard similarity of different contexts. Gen, Ret and Ctr respectively represent generated contexts, retrieved contexts and counter-memory contexts. Detailed discussion about Jaccard similarity is shown in Section 5.2.
@@
Following the above procedures, we obtain a subsets from original AIR datasets, with each instance encompassing a question $q$, a generated context $d_{1}^{\varrho}$, a retrieved context $d_{1}^{\gamma}$, a counter-memory context $d_{1}^{\varrho^{\prime}}$ and the associated answers for these contexts.
A.10.2 Discussion
In Table 3, we observe another phenomenon: GPT-3.5 exhibits a stronger inclination towards counter-memory contexts compared to the original generated contexts, which does not exist in Llama2-13b. This phenomenon is also found in the results of Xie et al. (2023). Upon further investigation, we discovered that this may be due to the relatively higher similarity betwee the counter-memory context generated by GPT-3.5 and the question, as shown in Table 4. We also analyse the conflictQA datasets proposed by Xie et al. (2023), and also find the similarity gap between generated and counter-memory contexts. This means that previous works may ignore some significant factors to LLMs’ preference, such as similarity (investigated in Section 5.2).
A.11 More Details and Results about Similarity
A.11.1 Similarity Metric
We employ Jaccard similarity to assess the term-based overlap, and BERTScore (Zhang et al., 2020) for evaluating the semantic similarity between contexts and questions. To mitigate the effect of length discrepancies between contexts and questions, we calculate the similarity at the sentence level and then aggregate them to derive the overall context-question similarity. In this section, we adopt a maximum aggregation strategy due to the single-hop nature of the NQ and TQA datasets, where the majority of questions can be answered using a small subset of sentences. We also try average aggregation strategy and observe similar results.
@@ A.11.2 Similarity Distribution
A.11.3 Effect of Similarity.
Figure 15 demonstrates a general trend that on slices with a smaller average similarity gap, LLMs exhibit a reduced preference for generated context. This trend is not very stable on NQ-AIR, which can be attributed to the smaller number of instances on NQ-AIR, as shown in Table 8.
Figure 13: BERTScore distribution of retrieved context and context generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
Figure 14: Jaccard Similarity distribution of retrieved context and context generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
(a) GPT 4
(b) GPT 3.5
(c) Llama2-13b
@@
A.11.4 Cases about Similarity
Table 12 shows examples that contain contexts with different similarity to the question. The contexts with high similarity typically directly support answering by repeating the phrasing in the question. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences. These observations indicate that text similarity can partly reflect the relevance between a question and a context, as well as the difficulty a LLM encounters in identifying potential answers.
TQA-AIR Example TQA-AIG Example Question Between 1959 and 1967 which city was the capital of Pakistan (Islamabad was being built)? Who is the most successful UK solo artist in the USA? Golden Answer Rawalpindi Elton John Generated Context Between 1959 and 1967, the capital of Pakistan was Karachi. Karachi is the largest city in Pakistan and is located on the southern coast of the country… Jaccard Similarity: 0.47 BertScore: 0.93 Elton John is the most successful UK solo artist in the USA. Born Reginald Kenneth Dwight in 1947, he adopted the stage name Elton John in the late 1960s… Jaccard Similarity: 0.69 BertScore: 0.93 Retrieved Context …was first shifted temporarily to Rawalpindi in the early 60s, and then to Islamabad when essential development work was completed in 1966… Jaccard Similarity: 0.16 BertScore: 0.85 …In 2009, Jay Sean’s single "Down" reached the number one spot on the "Billboard" Hot 100 and sold millions in the United States, making him ẗhe most successful male UK urban artist in US chart historyät the time… Jaccard Similarity: 0.14 BertScore: 0.86 Model output Karachi Elton John

### v4

@@ A.1.2 Retrieval-Augmented Approaches
The retrieval-augmented approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) enhance LLMs by incorporating relevant documents from the external corpus. These approaches represent a promising direction for addressing the knowledge limitations of LLMs, such as the need for knowledge updating (Jang et al., 2022) and long-tail knowledge (Kandpal et al., 2023). Early retrieval-augmented methods (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022) focused on the joint training of LLMs and retrieval modules to improve their cooperation. With the evolution of general-purpose LLMs, recent studies (Ram et al., 2023; Shi et al., 2023) investigate the strategy of appending relevant documents directly to the input while keeping the LLMs frozen. Despite the effectiveness of these methods, the retrieval-augmented approaches still face challenges due to irrelevant retrieval results and incomplete knowledge coverage Yu et al. (2023); Mallen et al. (2023). These noisy retrieval results can misguide the outcomes of LLMs (Mallen et al., 2023; Yoran et al., 2023; Ren et al., 2023).
A.1.3 Hybrid Approaches and Knowledge Conflicts
@@
Recent works investigate merging retrieved and generated contexts to leverage both parametric knowledge and external knowledge Abdallah and Jatowt (2023); Yu et al. (2022). These combination methods have shown improved performance over those relying solely on a single information source in a fully-supervised setting. Furthermore, (Zhang et al., 2023) proposes an improved method to effectively leverage the two sources of information, especially when conflicts arise. While these works have focused on improving the efficacy of hybrid approaches, the underlying mechanisms by which LLMs process conflicting information from different types of contexts remain underexplored.
@@
Current research on knowledge conflicts in LLMs primarily focuses on two aspects: conflicts between input contexts and LLMs’ parametric memory, and conflicts among the input contexts themselves. Regarding the former, Xie et al. (2023); Chen et al. (2022) find that LLMs are highly receptive to the input contexts rather than their internal memory. Concerning conflicts within multiple input contexts, Chen et al. (2022) demonstrates that LLMs tend to rely on a few most relevant retrieved contexts. Additionally, Xie et al. (2023) reveals a confirmation bias, i.e., LLMs demonstrate a tendency to favor contexts that align with their parametric knowledge when confronted with both supporting and opposing contexts. However, these studies are limited to analyzing context conflicts within a single type of input context. Our work complements these studies by considering conflicts between generated and retrieved contexts and reveals several key factors, such as semantic completeness and text similarity.
A.2 Length Control for Generated Contexts
A.2.1 Length Distribution across LLMs
@@
In our proposed framework, we regulate the length of generated contexts by incorporating length constraints in the prompt:
Generate a background context from Wikipedia to answer the given question {#question} Keep the length of the document around $n$ words
We observed that GPT 4 effectively controls the output length, whereas other models struggle with this aspect. To address this issue in the latter, we employ multiple values of $n$ and select the one that best matches the retrieved context.
As a result, Figure 9 shows the length distribution of retrieved contexts and contexts generated by various LLMs. The length distribution of retrieved contexts is more concentrated as they consist of text limited to precisely 100 words, along with their titles (Karpukhin et al., 2020). The variation in the length of different retrieved contexts is solely due to the differences in title lengths.
A.2.2 Length Distribution with Different Control Methods
In Section 5.3, we employ three methods, “Nature”, “Trunc.” and “S-Trunc.”, to vary the completeness of generated contexts, while controlling the length at the same time. Figure 10 illustrates the length distribution for generated contexts corresponding to these methods. From the results, we can observe that the contexts generated by original GenRead (Yu et al., 2022) are significantly longer compared to the retrieved contexts.
Table 8 illustrates the average similarity and completeness of three types of generated contexts. “Nature”, “Trunc.” and “S-Trunc.” result in contexts with similar average length and similarity, with preliminary differences in completeness.
Context Completeness NQ-AIR TQA-AIR Sentence Semantic Length Jaccard Length Jaccard GPT-3.5 Nature 106.0 0.187 105.0 0.398 S-Trunc 104.9 0.184 106.3 0.397 Trunc 107.4 0.187 103.6 0.397 Llama2-13b Nature 109.7 0.202 105.5 0.313 S-Trunc 105.9 0.193 104.8 0.293 Trunc 107.4 0.196 106.3 0.295
Table 8: Average length and similarity of generated context with different control methods. Three types of methods create contexts with comparable average length and similarity. The core difference lies in the completeness.
@@ A.3 Dataset Size
Table 9: The data quantities of the constructed subsets for different (Generator, Reader) pairs. NQ and TQA refer to the original datasets (dev+test).
A.4 Effect of Context Order
@@
generated-first 0.699 0.682
retrieved-first 0.665 0.556
random 0.691 0.586
Table 10: $\mathrm{DiffGR}$ with different context order on NQ-AIR and TQA-AIR datasets. GPT-3.5 serves as the generator and reader.
In the above experiments, retrieved and generated contexts are presented in random order. Previous studies Xie et al. (2023); Liu et al. (2023); Lu et al. (2022) have found that the model may be sensitive to the order of the input contexts. In their experiments, the input context was either all retrieved Liu et al. (2023) or all generated Xie et al. (2023). We conducted experiments to investigate whether the context order impacts the preference for the generated context. The generated and retrieved contexts are concatenated with three different orders: generated-first, retrieved-first, and random. To control the cost of API, this section conducts experiments on the context-conflicting datasets from only the test sets of NQ and TQA. We compute the $\mathrm{DiffGR}$ with different context orders respectively.
As shown in Table 10, across all context orders, LLMs consistently show a strong tendency to favor generated contexts. When the retrieved context is positioned first, there is a slight reduction in $\mathrm{DiffGR}$. This reduction may result from the LLMs’ preference for generated contexts being partially offset by their bias towards the top context (Liu et al., 2023; Xie et al., 2023).
A.5 More Results on AIG Datasets
Figure 11 shows the $\mathrm{DiffGR}$ with different (reader, generator) pairs on their corresponding NQ-AIG datasets. It can be observed that LLMs show a strong tendency to rely on generated contexts across various (reader, generator) pairs.
Figure 11: $\mathrm{DiffGR}$ with different (reader, generator) pairs on their corresponding NQ-AIG datasets.
A.6 More Details about Confirmation Bias
A.6.1 Experiment Setting
We construct counter-memory contexts for each instance on the original AIR context-conflicting subsets, as outlined below:
@@ A.6.1 Experiment Setting
Give you a reference word, transform it into an analogous word.
Reference: Missouri River
Analogous Word: Mississippi River
Reference:
{memory answer}
Analogous Word:
Furthermore, to facilitate the calculation of $\mathrm{DiffGR}$, we only retain those samples where the counter-memory answer also diverges from $a_{\phi}^{\gamma}$, the answer provided by the retrieved context. Formally, the counter-memory answer satisfies $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\varrho}$ and $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\gamma}$.
@@
Answer Consistency Checking. To verify that the counter-memory context actually supports the counter-memory answer, we retain only those instances where the predicted answer, derived exclusively from the counter-memory context, exactly matches the counter-memory answer:
@@
Context NQ-AIR TQA-AIR ConflictQA(PopQA) Length Jaccard Length Jaccard Length Jaccard Retrieved 107.5 0.111 106.3 0.184 - - GPT-3.5 Gen 105.9 0.194 104.9 0.385 62.14 0.169 Ctr 105.7 0.251 103.8 0.499 100.4 0.236 Llama2-13b Gen 107.4 0.197 104.3 0.319 - - Ctr 106.7 0.165 103.5 0.310 - -
Table 11: Average length and Jaccard similarity of different contexts. Gen, Ret, and Ctr respectively represent generated contexts, retrieved contexts, and counter-memory contexts. “-” means the context is not included in this dataset. Detailed discussion about Jaccard similarity is shown in Section 5.2.
@@
Following the above procedures, we obtain subsets from original AIR datasets, with each instance encompassing a question $q$, a generated context $d_{1}^{\varrho}$, a retrieved context $d_{1}^{\gamma}$, a counter-memory context $d_{1}^{\varrho^{\prime}}$ and the associated answers for these contexts. Table 12 shows the size of developed context-conflicting subsets with counter-memory contexts.
A.6.2 Discussion
In Table 4, we observe a phenomenon: GPT-3.5 exhibits a stronger bias towards counter-memory contexts compared to the original generated contexts, which does not exist in Llama2-13b. This phenomenon is also observed in Xie et al. (2023) (Table 5 in their work). Upon further investigation, we discover that this may be due to the relatively higher similarity between the counter-memory context generated by GPT-3.5 and the question, as shown in Table 11. In contrast, the generated context and counter-memory context for Llama2-13b exhibit approximately equal levels of similarity. Furthermore, we analyze the original ConflictQA (PopQA) datasets introduced by Xie et al. (2023) and identify a similar disparity in context similarities, as shown in Table 11. This means that previous works may ignore some significant factors to LLMs’ preference, such as similarity (investigated in Section 5.2).
A.7 More Details and Results about Similarity
A.7.1 Similarity Metric
We employ Jaccard similarity to assess the term-based overlap, and BERTScore (Zhang et al., 2020) for evaluating the semantic similarity between contexts and questions. To mitigate the effect of length discrepancies between contexts and questions, we calculate the similarity at the sentence level and then aggregate them to derive the overall context-question similarity. In this work, we adopt a maximum aggregation strategy due to the single-hop nature of the NQ and TQA datasets, where the majority of questions can be answered using a small subset of sentences. We also try the average aggregation strategy and observe similar results.
@@ A.7.2 Similarity Distribution
A.7.3 Effect of Similarity.
Figure 15 demonstrates a general trend that on slices with a smaller average similarity gap, LLMs exhibit a reduced preference for generated context.
Figure 13: BERTScore distribution of retrieved contexts and contexts generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
Figure 14: Jaccard Similarity distribution of retrieved contexts and contexts generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
@@
A.7.4 Cases about Similarity
Table 13 shows examples that contain contexts with different similarities to the question. The contexts with high similarity typically directly support answering by repeating the phrasing in the question. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences. These observations indicate that text similarity can partly reflect the relevance between a question and a context, as well as the difficulty the LLM encounters in identifying potential answers.
TQA-AIR Example TQA-AIG Example Question Between 1959 and 1967 which city was the capital of Pakistan (Islamabad was being built)? Who is the most successful UK solo artist in the USA? Golden Answer Rawalpindi Elton John Generated Context Between 1959 and 1967, the capital of Pakistan was Karachi. Karachi is the largest city in Pakistan and is located on the southern coast of the country… Elton John is the most successful UK solo artist in the USA. Born Reginald Kenneth Dwight in 1947, he adopted the stage name Elton John in the late 1960s… Jaccard Similarity: 0.47 Jaccard Similarity: 0.69 BertScore: 0.93 BertScore: 0.93 Retrieved Context …was first shifted temporarily to Rawalpindi in the early 60s, and then to Islamabad when essential development work was completed in 1966… …In 2009, Jay Sean’s single "Down" reached the number one spot on the "Billboard" Hot 100 and sold millions in the United States, making him ẗhe most successful male UK urban artist in US chart historyät the time… Jaccard Similarity: 0.16 Jaccard Similarity: 0.14 BertScore: 0.85 BertScore: 0.86 Model output Karachi Elton John
@@ A.7.4 Cases about Similarity
A.8 Cases about completeness
Table 14 provides some examples to facilitate the understanding of completeness. From the cases, we observe that retrieved contexts and “Trunc.” often contain incomplete sentences. Additionally, compared to “S-Trunc”, “Nature” typically exhibits greater semantic completeness. Specifically, “Nature” often encompasses a full logical structure of an article, including an introduction, discussion, and conclusion, whereas “S-Trunc” may terminate abruptly.
Example Question: In World War I, what did the Allies call the defence line that was known as the ’Siegfried Line’ to the Germans? Answer: [’Hindenberg Line’, ’Hindenburg line’, ’Battles of the Hindenburg Line’, "’HINDENBURG LINE’", ’Advance to the Hindenburg Line’, ’Hindenburg Line’] Retrieved context: Title: Siegfried Line Content: Siegfried Line The term Siegfried Line refers to two different German defensive lines, one during the First World War and the other during the Second World War. The original was a line of forts and tank defences built by Germany in northern France during 1916–17, known in German as the Siegfriedstellung ("Siegfried Position"). The British referred to it as the "Hindenburg Line", a name subsequently adopted by the rest of the Allies. After WW1 the German Fortifications East of the Rhein were required to be demolished due to the Treaty of Versailles. The second, commonly known in German as the Generated context (Nature): The Allies referred to the German defence line known as the ’Siegfried Line’ as the ’Westwall’ during World War I. Constructed between 1936 and 1940, the Westwall was a system of fortifications built along Germany’s western border. It stretched from the Swiss border to the Dutch border, covering approximately 630 kilometers. The line consisted of concrete bunkers, tank traps, and anti-tank ditches, aiming to deter any potential invasion from the west. The Allies recognized the formidable nature of this defensive barrier and thus referred to it as the Westwall, while the Germans named it the Siegfried Line, after the mythical hero Siegfried from German folklore. Generated context (Trunc.): The Allies referred to the German defence line known as the "Siegfried Line" as the "Westwall" during World War I. The Westwall was a series of fortifications constructed by Germany along its western border, stretching from the Swiss border to the Netherlands. It was named after the German Emperor, Wilhelm II, who was commonly referred to as "Kaiser Wilhelm" or "Siegfried" by the Allies. The construction of the Westwall began in the 1930s as a response to the growing tensions in Europe and the threat of war. The line consisted of a network of bunkers, pillboxes, tank traps, and other defensive structures, strategically positioned Generated context (S-Trunc.): The Allies referred to the German defence line known as the "Siegfried Line" as the "Westwall" during World War I. The Westwall was a series of fortifications constructed by Germany along its western border, stretching from the Swiss border to the Netherlands. It was named after the German Emperor, Wilhelm II, who was commonly referred to as "Kaiser Wilhelm" or "Siegfried" by the Allies. The construction of the Westwall began in the 1930s as a response to the growing tensions in Europe and the threat of war. The line consisted of a network of bunkers, pillboxes, tank traps, and other defensive structures, strategically positioned to impede any potential invasion from the west.
Table 14: Examples with retrieved contexts and generated contexts. “Nature”, “Trunc.” and “S-Trunc.” represent three types of generated contexts with different completeness. Retrieved contexts often contain incomplete sentences.
Reader Generator NQ-AIG NQ-AIR TQA-AIG TQA-AIR Gen Ret Others Gen Ret Others Gen Ret Others Gen Ret Others GPT 4 GPT 4 0.9125 0.0589 0.0286 0.7379 0.1743 0.0878 0.9387 0.0304 0.031 0.7651 0.1762 0.0587 GPT 3.5 GPT 3.5 0.9177 0.0449 0.0374 0.7083 0.1470 0.1447 0.9347 0.026 0.0393 0.7332 0.1775 0.0893 Llama2-13b Llama2-13b 0.8966 0.0500 0.0534 0.7216 0.1811 0.0973 0.9071 0.0433 0.0496 0.7212 0.1918 0.0870 Llama2-7b Llama2-7b 0.7041 0.1082 0.1876 0.6148 0.2071 0.1781 0.7973 0.0927 0.1100 0.6555 0.2101 0.1344
Table 15: “Gen” denotes the proportion of responses that match the candidate answer within generated contexts, whereas “Ret” refers to the proportion of matching the candidate answer within retrieved contexts. “Others” encompasses the proportion of responses that do not align with either category. Given that the proportion of “Others” is significantly lower relative to the disparities between “Gen” and “Ret”, its impact on the conclusions of this paper is negligible.